# OPERA: Online Data Pruning for Efficient Retrieval Model Adaptation

### 17 Mar 2026 v1

@@ 1 Introduction
Data pruning and coreset selection have shown promise for improving training efficiency in neural networks (Toneva et al., 2018; Marion et al., 2023; Sorscher et al., 2022; Killamsetty et al., 2021), and dynamic pruning methods have further demonstrated that adjusting data selection during training can maintain performance while reducing computation (Qin et al., 2023; Li et al., 2024; Huang et al., 2024). However, these methods are designed for standard classification or generation tasks where training samples are treated as independent, identically distributed instances. Dense retriever finetuning is fundamentally different: it employs a two-stage contrastive sampling framework (Xiao et al., 2023; Chen et al., 2024; Zhang et al., 2023) where queries are first sampled, and then positive and negative documents are selected for each query. This hierarchical structure means that data quality operates at two distinct granularities (query relevance and document relevance), creating unique challenges that existing pruning methods do not address. To our knowledge, no prior work has studied data pruning specifically for dense retriever finetuning. A detailed discussion of related work is provided in Appendix A.
@@
  • We validate OPERA across 8 datasets, 6 domains, and 2 architectures (encoder-only BGE and decoder-based Qwen3-Embedding), with theoretical guarantees on when pruning outperforms standard finetuning.
@@ 2.1 Preliminary: Standard Finetuning (FT)
Given a dataset $Q$ with $n$ queries where each query $q\in Q$ has $m_{q}$ positive documents, we adopt a contrastive learning framework with two-stage sampling (Xiao et al., 2023). For each training step, we first uniformly sample queries from $Q$. Then, for each sampled query, we randomly select one positive document $d$ from its $m_{q}$ positive documents and one negative document through hard negative mining for loss computation.
@@
$P_{t}(q)=\nicefrac{{1}}{{n}},\quad P_{t}(d|q)=\nicefrac{{1}}{{m_{q}}}$ (1)
Note that for notational simplicity, we omit the time step $t$ for the following equations.
@@ 2.2 Static Pruning (SP)
To understand how data quality affects retrieval finetuning, we begin with a straightforward approach inspired by Sorscher et al. (2022): retaining only the highest-quality training pairs. We compute cosine similarity between each query-document pair using the pretrained model and retain only the top fraction of pairs for training. Mathematically, we use an indicator function $I$ to mark whether a query or query-document pair is kept:
$I=\begin{cases}1,&\text{if kept}\\ 0,&\text{if pruned}\end{cases}$ (2)
@@
$P(q)=\frac{\sum_{d}I_{d}}{\sum_{q}\sum_{d}I_{d}},\quad P(d|q)=\frac{I_{d}}{\sum I_{d}}$ (3)
In practice, we select top $k\sum m_{q}$ query-document pairs with highest similarity scores for training, where $k$ is the data retention rate:
$k=\frac{\sum_{q}\sum_{d}I_{d}}{\sum m_{q}}$ (4)
@@ 2.3 Dynamic Pruning (DP)
The sampling process, detailed in Algorithm 1, consists of two components. For query sampling, we combine top-scoring queries with randomly selected low-scoring queries to maintain diversity. In document sampling, we assign sampling weights to documents based on their quality while ensuring all documents maintain a non-zero probability of selection.
@@
Update Interval To mitigate the computational overhead associated with frequent score updates and pruning operations, we introduce an update interval parameter $I_{u}$. We update the query scores $\mathbf{q}$ every $I_{u}$ iterations, while maintaining the per-iteration random selection of $d$ due to its negligible computational cost. Our empirical analysis demonstrates that this optimization reduces additional computation time from 4.5% to 1.64% over the baseline, without impacting performance. The effects of varying $I_{u}$ are examined in Section 3.4.2.
2.4 Theoretical Analysis: When Does Pruning Help?
Lemma 1.
Proof.
Direct computation yields $f^{\prime}(k)=\frac{(1-(u^{T}v)^{2})(1-k)}{\|ku+(1-k)v\|^{3}}.$ Since $u\neq v\implies u^{T}v<1$, and $k<1$, thus $f^{\prime}(k)>0$. ∎
Theorem 1.
Then:
$E^{SP}>E^{FT}\Leftrightarrow\gamma>\frac{m_{q}^{+}}{m_{q}}$ (7)
and similarly for DP. Additionally, when $\gamma=\rho$:
$E^{SP}>E^{DP}.$ (8)
Proof.
and similarly for SP and DP:
$\quad+((1\!-\!\rho)s(\beta\!-\!1)+m_{q}\!-\!m_{q}^{+})\mu_{2})$ (10)
where $C_{1},C_{2},C_{3}$ are normalization constants. By Lemma 1:
and similarly for DP. Comparing SP and DP when $\gamma=\rho$:
$E^{SP}>E^{DP}\Leftrightarrow\gamma>\tfrac{m_{q}^{+}}{m_{q}}$ (12)
Theorem 1 provides a general justification for quality-based pruning: any scoring function that identifies true positives at a rate higher than the dataset’s base rate ($m_{q}^{+}/m_{q}$) will improve learned query representations. Under equal selection quality, SP outperforms DP because it completely excludes noisy samples rather than merely down-weighting them. However, as training progresses, DP can surpass SP through improved sampling quality ($\rho$) and strength ($\beta$) enabled by dynamic threshold scheduling. This motivates a two-stage approach: applying SP first to discard strong noise, followed by DP to refine on the filtered data. The empirical validation is presented in Section 3.5.
@@ 3 Experiments
  • RQ6: How does DP’s sampling behavior evolve during training?
@@ 3.2 Main Results
Table 2: OPERA vs. baselines on Qwen3-Embedding-0.6B. Best in bold, second-best underlined.
@@
SP confirms the quality-coverage tradeoff in the previous section: it outperforms FT and InfoBatch in NDCG@10 across unseen, seen, and all datasets on average (average rank: 2.63), achieving the best NDCG@10 on NFCorpus (0.491), TriviaQA (0.501), and FEVER (0.915). As expected from our analysis, this comes at the cost of recall, as SP’s Recall@20 average rank drops to 3.88, because a priori removal of training pairs reduces query diversity. Despite this, SP offers substantial data efficiency: it drops 75% of pairs yet improves ranking, and subsequent analysis demonstrates even faster convergence and effective denoising capabilities.
@@ 3.6 Sampling Weight Visualization
To analyze how DP allocates training focus over time, we visualize the evolution of sampling probabilities on the FiQA (Maia et al., 2018) dataset (Figure 2). Unlike curriculum learning (Bengio et al., 2009), which imposes a fixed easy-to-hard ordering and may exclude examples at certain stages, DP keeps all examples accessible throughout training while continuously re-evaluating and adjusting their sampling weights as model representations evolve.
Figure 2(a) shows the probability distribution grouped by query. Crucially, nearly all queries maintain nonzero sampling probabilities throughout training, demonstrating that DP preserves broad query coverage unlike SP, which would exclude many queries entirely. The variation in intensity reflects quality-aware up-weighting while maintaining the diversity needed for recall. Figure 2(b), sorted by initial probabilities, reveals how DP dynamically redistributes attention: initially high-probability examples may decrease in importance while previously low-priority examples gain prominence. This demonstrates DP’s ability to discover valuable training examples beyond what the pretrained model initially favors, shifting focus toward examples more informative for domain-specific adaptation.
@@
Figure 2: DP sampling probability evolution. Color: black (low) to yellow (high).
@@ A.2 Data Pruning in Neural Networks
A.3 Curriculum Learning
Curriculum learning has emerged as a promising approach to train neural networks by presenting training examples in a meaningful order. The idea of curriculum learning was formalized in (Bengio et al., 2009), where it showed that gradually increasing the difficulty of training examples could lead to better generalization and faster convergence. In natural language processing, CL has been applied to machine translation (Platanios et al., 2019), question answering (Sachan and Xing, 2016), language modeling (Graves et al., 2017), etc. Several approaches have been proposed for automatically determining the difficulty of training examples and generating curricula. Self-paced learning (Kumar et al., 2010) allows the model to automatically select its own curriculum based on the loss of training examples. Other works have explored the use of teacher-student frameworks (Matiisen et al., 2019), where a teacher model determines the curriculum for a student model.
Our dynamic pruning framework incorporates curriculum learning principles through its evolving threshold scheduling, which gradually adjusts both the ratio and sampling strength of training instances as the model’s representations become more refined. However, a key distinction is that OPERA selects examples based on data quality (query-document similarity) rather than difficulty. These are distinct axes: a high-similarity query-document pair can be easy or hard depending on the negative mining strategy, and a difficult example is not necessarily high-quality. This quality-based perspective, combined with the hierarchical query-document sampling structure specific to retrieval, differentiates our approach from standard curriculum learning.
@@ Appendix B Theoretical Analysis
The full theoretical analysis, including Lemma 1, Theorem 1, and their proofs, is presented in Section 2.4 of the main paper.

### 01 Apr 2026 v2

@@ 1 Introduction
Data pruning and coreset selection have shown promises for improving training efficiency in neural networks (Toneva et al., 2018; Marion et al., 2023; Sorscher et al., 2022; Killamsetty et al., 2021), while dynamic pruning methods have particularly demonstrated that adjusting data selection during training can maintain performance while reducing computation (Qin et al., 2023; Li et al., 2024; Huang et al., 2024). However, these methods are designed for standard classification or generation tasks where training samples are treated as independent, identically distributed instances. Dense retriever finetuning is fundamentally different: it employs a two-stage contrastive sampling framework (Xiao et al., 2023; Chen et al., 2024; Zhang et al., 2023) where queries are first sampled, and then positive and negative documents are selected for each query. This hierarchical structure means that data quality operates at two distinct granularities (query relevance and document relevance), creating unique challenges that existing pruning methods do not address. To our knowledge, no prior work has studied data pruning specifically for dense retriever finetuning. A detailed discussion of related work is provided in Appendix A.
Refer to captionRefer to caption
Figure 1: Comparison of sampling probability distributions across training strategies. Left three panels: Standard finetuning (FT) samples all data pairs uniformly, while static pruning (SP) discards the lowest-similarity ones and up-weights the rest, improving ranking but reducing query coverage. Dynamic pruning (DP) at the start of training retains all pairs with non-zero probability, assigning higher sampling weights to high-similarity pairs through soft thresholding. As training progresses, DP redistributes sampling weights: the model discovers informative examples in the low/mid similarity range and up-weights them. Right panel: Average rank of each method on NDCG@10 and Recall@20 across 8 datasets (lower rank is better); DP achieves the best rank on both metrics. See Section 3.6 for empirical sampling weight evolution on real data.
@@
  • We validate OPERA across 8 datasets, 6 domains, and 2 architectures (encoder-only BGE and decoder-based Qwen3-Embedding). We further demonstrate effective denoising through a combined SP+DP strategy, and provide theoretical guarantees on when pruning outperforms standard finetuning.
@@ 2.1 Preliminary: Standard Finetuning (FT)
Given a dataset $(Q,D)$ with $n$ queries where each query $q\in Q$ has $m_{q}$ positive documents from the document corpus $D$, we adopt a contrastive learning framework with two-stage sampling (Xiao et al., 2023). For each training step $t$, we first uniformly sample queries from $Q$. Then, for each sampled query, we randomly select one positive document $d$ from its $m_{q}$ positive documents and one negative document $d^{\prime}$ through hard negative mining (HNM) for contrastive loss computation (Xiao et al., 2023).
@@
$P_{t}(q)=\nicefrac{{1}}{{n}},\quad P_{t}(d|q)=\nicefrac{{1}}{{m_{q}}},\quad P_{t}(d^{\prime}|q)=HNM(D,q)$ (1)
For notational simplicity, we omit the time step $t$ and $P(d^{\prime}|q)$ for the following equations.
@@ 2.2 Static Pruning (SP)
To understand how data quality affects retrieval finetuning, we begin with a straightforward approach inspired by Sorscher et al. (2022): retaining only the highest-quality training pairs. We compute cosine similarity between each query-document pair using the pretrained model and retain only the top fraction of pairs for training. We define $D(q)$ as the set of positive documents for query $q$, and use indicator variables $I_{q}$ and $I_{q,d}$ to denote whether a query or query-document pair is kept by the SP algorithm:
$I_{q},I_{q,d}=\begin{cases}1,&\text{if kept}\\ 0,&\text{if pruned}\end{cases}$ (2)
@@
$P(q)=\frac{\sum_{d\in D(q)}I_{q,d}}{\sum_{q\in Q}\sum_{d\in D(q)}I_{q,d}},\quad P(d|q)=\frac{I_{q,d}}{\sum_{d\in D(q)}I_{q,d}},$ (3)
where $Q$ is the set of all queries. In practice, we select top $k\sum_{q\in Q}m_{q}$ query-document pairs with highest similarity scores for training, where $k\in[0,1]$ is the data retention rate:
$k=\frac{\sum_{q\in Q}\sum_{d\in D(q)}I_{q,d}}{\sum_{q\in Q}m_{q}}$ (4)
@@ 2.3 Dynamic Pruning (DP)
As detailed in Algorithm 1, DP uses two scores updated from the model’s forward pass at each training step: the positive score $\mathbf{p}_{i}=s(e_{q},e_{d_{i}})$, the cosine similarity between query and positive document embeddings, and the query score $\mathbf{q}_{j}$, the mean contrastive loss across query $j$’s positive pairs (offset to ensure positivity). These drive a two-level sampling process: at the query level, top-scoring queries are combined with randomly selected low-scoring queries to maintain diversity; at the document level, sampling weights are assigned based on positive scores while ensuring all documents retain non-zero selection probability.
@@
Update Interval. To mitigate the computational overhead associated with frequent score updates and pruning operations, we introduce an update interval parameter $I_{u}$. We update the query scores $\mathbf{q}$ every $I_{u}$ iterations, while maintaining the per-iteration random selection of $d$ due to its negligible computational cost. Our empirical analysis demonstrates that this optimization reduces additional computation time from 4.5% to 1.64% over the baseline, without impacting performance. The effects of varying $I_{u}$ are examined in Section 3.4.2.
Comparison with Adaptive Training Methods. Table 1 compares DP with related adaptive training paradigms (see Appendix A for a broader discussion). Curriculum learning (Bengio et al., 2009), SPL (Kumar et al., 2010), and active learning all rely on scores computed before or between training stages rather than continuous scoring from the current model weights, and permanently exclude samples rather than maintaining full coverage. Additionally, curriculum learning follows a predetermined ordering that is not adaptive to model state, while SPL and active learning require expensive scoring or labeling passes. InfoBatch addresses these limitations, but applies uniform sample-level pruning without hierarchical sampling or scheduled strength. DP performs lightweight adaptive online update with continuous scoring and full data coverage, uses a scheduled sampling strength, and operates at two hierarchical levels reflecting the query-document structure of retrieval training.
Lightweight Adaptive Continuous Scoring Full Coverage Hierarchical Scheduled
Curriculum
SPL
Active Learning
InfoBatch
DP (ours)
Table 1: Comparison of adaptive training paradigms.
When Does Pruning Help? We provide a formal analysis of the conditions under which data pruning outperforms standard finetuning (Appendix B). The key result (Theorem 1) shows that any scoring function identifying true positives at a rate higher than the noise rate will improve learned query representations. Under equal selection quality, SP outperforms DP because it completely excludes noisy samples rather than merely down-weighting them. However, when signal quality is low, SP’s hard removal reduces the training pool while retaining noise, increasing per-sample exposure to mislabeled examples and the risk of overfitting. DP avoids this by keeping all samples accessible with non-zero probability and increasing its sampling strength $\alpha$ over training as the model improves, becoming more selective when the signal is more reliable. This may explain why DP outperforms SP when initial signal is not robust. This motivates a two-stage approach: applying SP first to discard strong noise, followed by DP to refine. The empirical validation is presented in Section 3.5.
@@ 3 Experiments
  • RQ6: How does DP allocate training focus across examples over time?
@@ 3.2 Main Results
Table 3: Comparison on Qwen3-Embedding-0.6B. Best in bold, second-best underlined.
@@
SP confirms the quality-coverage tradeoff in the previous section: it outperforms FT and InfoBatch in NDCG@10 across unseen, seen, and all datasets on average, achieving the best NDCG@10 on NFCorpus (0.491), TriviaQA (0.501), and FEVER (0.915). As expected from our analysis, this comes at the cost of recall, as SP’s Recall@20 average rank drops to 3.88, because a priori removal of training pairs reduces query diversity. Despite this, SP offers substantial data efficiency: it drops 75% of pairs yet improves ranking, and subsequent analysis demonstrates even faster convergence and effective denoising capabilities.
@@ 3.6 Sampling Weight Visualization
To analyze how DP allocates training focus over time, we visualize the evolution of sampling probabilities on the FiQA (Maia et al., 2018) dataset (Figure 3). Like self-paced learning (Kumar et al., 2010), DP continuously re-evaluates and adjusts sampling weights as model representations evolve, while ensuring all examples remain accessible throughout training (see Section 1 for the detail comparison).
Figure 3(a) shows the probability distribution grouped by query. Notably, nearly all queries maintain nonzero sampling probabilities throughout training, demonstrating that DP preserves broad query coverage unlike SP, which would exclude many queries entirely. The variation in intensity reflects quality-aware up-weighting while maintaining the diversity needed to preserve and even improve recall. Figure 3(b), sorted by initial probabilities, reveals how DP dynamically redistributes attention: initially high-probability examples may decrease in importance while previously low-priority examples gain prominence. This demonstrates DP’s ability to discover valuable training examples beyond what the pretrained model initially favors, shifting focus toward examples more informative in a domain-specific manner during the model training.
@@
Figure 3: DP sampling probability evolution.
@@ A.2 Data Pruning in Neural Networks
A.3 Curriculum Learning and Adaptive Training
Curriculum learning has emerged as a promising approach to train neural networks by presenting training examples in a meaningful order. The idea of curriculum learning was formalized in (Bengio et al., 2009), where it showed that gradually increasing the difficulty of training examples could lead to better generalization and faster convergence. In natural language processing, CL has been applied to machine translation (Platanios et al., 2019), question answering (Sachan and Xing, 2016), language modeling (Graves et al., 2017), etc. Several approaches have been proposed for automatically determining the difficulty of training examples and generating curricula. Self-paced learning (SPL) (Kumar et al., 2010) allows the model to automatically select its own curriculum based on the loss of training examples. Other works have explored the use of teacher-student frameworks (Matiisen et al., 2019), where a teacher model determines the curriculum for a student model. Related paradigms include active learning, which selects informative samples in discrete select-retrain cycles (Ren et al., 2021), and self-training, which iteratively refines pseudo-labels (Xie et al., 2020). InfoBatch (Qin et al., 2023) downsamples low-loss examples with inverse-probability reweighting.
Our dynamic pruning framework incorporates curriculum learning principles through its evolving threshold scheduling, which gradually adjusts both the ratio and sampling strength of training instances as the model’s representations become more refined. However, DP differs from all of the above in three key respects: (1) it operates at two hierarchical levels (query and document), reflecting the two-stage sampling structure specific to retrieval training; (2) it uses increasing sampling strength via cosine scheduling, becoming more selective as the model’s scoring improves, rather than relying on a fixed threshold or predetermined ordering; and (3) it maintains non-zero sampling probability for all examples throughout training, preserving coverage unlike curriculum learning and active learning which exclude examples entirely. A detailed comparison is provided in Section 1.
@@ Appendix B Theoretical Analysis
Lemma 1.
Proof. Direct computation yields $f^{\prime}(k)=\frac{(1-(u^{T}v)^{2})(1-k)}{\|ku+(1-k)v\|^{3}}.$ Since $u\neq v\implies u^{T}v<1$, and $k<1$, thus $f^{\prime}(k)>0$. ∎
Theorem 1.
Then:
$E^{SP}>E^{FT}\Leftrightarrow\gamma>\frac{m_{q}^{+}}{m_{q}}$ (7)
and similarly for DP. Additionally, when $\gamma=\rho$:
$E^{SP}>E^{DP}.$ (8)
Proof.
and similarly for SP and DP:
$\quad+((1\!-\!\rho)s(\beta\!-\!1)+m_{q}\!-\!m_{q}^{+})\mu_{2})$ (10)
where $C_{1},C_{2},C_{3}$ are normalization constants. By Lemma 1:
and similarly for DP. Comparing SP and DP when $\gamma=\rho$:
$E^{SP}>E^{DP}\Leftrightarrow\gamma>\tfrac{m_{q}^{+}}{m_{q}}$ (12)